OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. P(6, 2) Q(8, -4) I don't get it!!!

OpenStudy (anonymous):

You should recognize three form of a straight line: Standard Form: Ax + By = C Slope-Intercept Form: y = mx + b Point-Slope Form: y - y(sub1) = m (x - x(sub1)) Now find the slope m = (2 + 4)/(6 - 8) = 6/-2 = -3 Next, you can use either slope-intercept form or point slope form. I will use the point slope form and the point (6, 2) y - 2 = -3 (x - 6) = -3x +18 3x + y = 20